silvernight269:

Need help 2x^2+10x-14=0

silvernight269:

I tried 2(x^2+5x-7)=0

HanAkoSolo:

How are you trying to solve this? are we using the quadratic formula, are we solving for x, are we factoring? @silvernight269

silvernight269:

Solving for x

HanAkoSolo:

Okay. Honestly, this can be done with the quadratic formula

HanAkoSolo:

Just for review, it looks like this: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

silvernight269:

To make it equal to 0?

jhonyy9:

just one moment ? x_1,2 = this formule bc. there are + and - so result two roots

HanAkoSolo:

there are, but the roots are not exact

HanAkoSolo:

\(ax^2 + bx = c\) so we have to find the a, b, and c terms and then we can plug those in and solve for the various x's

HanAkoSolo:

So this means that a would equal 2, b =10, and c would be 14 since when you bring it to the other side, it becomes positive 14.

silvernight269:

It still doesn’t make sense

HanAkoSolo:

okay. Let's try plugging everything in then

HanAkoSolo:

\(x_{(1,2)} = \frac{ -10 \pm \sqrt{(10)^2 - 4(2)(14)} }{ 2(2) }\)

silvernight269:



silvernight269:



HanAkoSolo:

It would actually look more along the lines of this: x_{1,2} = \frac{ -10 \pm \sqrt{ 212 } }{ 4 }

HanAkoSolo:

oh, oops. Let's try that again: \(x_{1,2} = \frac{ -10 \pm \sqrt{ 212 } }{ 4 }\)

silvernight269:

Why 212

HanAkoSolo:

because I'm realizing that I goofed in telling you that c = 14. It's actually -14. The reason for this is because the equation you gave was already in quadratic form and I didn't realize. It's the negative that changes the answer so drastically \(10^2 - 4 (2) (-14)\) = 212

silvernight269:

Oooooooh

HanAkoSolo:

\(x_1 = \frac{ -10~+~\sqrt{ 212 } }{ 4 } = -\frac{5}{2}+\frac{1}{2}\sqrt{53}\) \(x_2 = \frac{ - 10~-~\sqrt{ 212 } }{ 4} = -\frac{5}{2} - \frac{1}{2}\sqrt{53}\)

HanAkoSolo:

This is what things should look like once everything is simplified, including the radicals. Does this make sense?

silvernight269:

Kinda

HanAkoSolo:

What do you still have questions on?

HanAkoSolo:

Like, you know, where in the process?

silvernight269:

Actually I think I’m ok

HanAkoSolo:

you sure?

silvernight269:

Yea

HanAkoSolo:

Okay, glad to have helped!